Circular No. 7354

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/2000 B2 (LINEAR) An object with unusual motion and reported as asteroidal by LINEAR

(discovery observation below) has been found to be cometary in appearance following posting on the NEO Confirmation Page. CCD observations by P. Kušnirák (Ondřejov, 0.65-m f/3.6 reflector) and by M. Tichý and Z. Moravec (Kleť, 0.57-m f/5.2 reflector) indicate that the object appears slightly diffuse. Complete astrometry and parabolic orbital elements (T = 1999 Nov. 10.5 TT, q = 3.78 AU) appear on MPEC 2000-C04.

2000 UT  α_{2000}  δ_{2000}  m_{1}
Jan. 29.23907  6^{h}38^{m}59.^{s}78  +36^{o}32′39′′.9  18.8

SUPERNOVA 1987A IN THE LARGE MAGELLANIC CLOUD P. Bouchet, Cerro Tololo Interamerican Observatory (CTIO); S. Law-

rence, A. Crotts, B. Sugerman, and R. Uglesich, Columbia University; and S. Heathcote, CTIO, report brightening of the ‘hotspot’ (IAUC 6665, 6710, 6761, 7056) and the probable appearance of new structure on the inner ring around SN 1987A: “Near-infrared images obtained on 1999 Dec. 25 using the CTIO 4-m telescope (+ Tip-Tilt system and OSIRIS) in a narrow band centered on He I 1083.0-nm show that the hotspot has brightened by a factor of 1.4 since 1998 Oct. 6 (cf. IAUC 7056; note that a different method for defining aperture photometry is used now). We also observe the probable appearance of a new source, perhaps an extension to the hotspot but consistent with a new point source at a radius of 0′′.6 and p.a. 104^{o} with respect to the supernova, near the inner ring edge on the side farther from the earth, and corresponding to a strong, previously known, inward protrusion of the ring. The original hotspot now contains ∼ 21 percent of the entire He I 1083.0-nm flux seen in the ring (including hotspots), and the new source contains ∼ 1 percent of the total ring flux. Meanwhile, the western half of the ring (with no apparent hotspots) has faded by 4 percent. This new source is the most significant sign of new activity beyond the first hotspot that we have seen in the course of our intensive monitoring program, and it may presage the appearance of structure at other wavelengths due to the ring/ejecta interaction.”

2000 February 1  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green